OpenStudy (anonymous):

A right triangle has a hypotenuse of length 15 and a leg of length 9. What is the length of the other leg of the triangle?

OpenStudy (futureescape):

a sqaured + b squared = c squared 9^2+b^2= 15^2 81+b^2=225 b^2= 144 square root of 144 =12 b^2 = 12^2 The length of the other leg is 12 squared. Hope this helps.

OpenStudy (anonymous):

thank you it helps out a lot